# Large Language Models Can Learn Temporal Reasoning

### 12 Jan 2024 v1

@@ 3.2.2 Data Augmentation
Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023). To address the issue, we propose two special strategies for data augmentation.
Learning with Disturbances   Essentially, our framework involves two steps: text-to-tempgraph translation and symbolic temporal reasoning. At the test stage, reasoning happens on the noisy graphs, which can impact the overall performance. Specifically, for the reasoning module, we train with ground truth/verified graphs while testing on the estimations. To improve the robustness of reasoning, we introduce some disturbances on the graphs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the model. We investigate two types of disturbances: remove irrelevant edges or replace edges by changing relation names, e.g., "owned" $\rightarrow$ "run", "was affiliated to" $\rightarrow$ "belonged to", "played for" $\rightarrow$ "joined". An edge is considered as irrelevant if it is not involved in both QA and CoTs.
@@
TempGraph (Adding Disturbances):
@@
Graph-based QA (Inductive Transformation):
@@
Table 3: We propose two special data augmentation strategies (adding disturbances and inductive transformation) for the neuro-symbolic reasoning.
where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation name changing $f_{R}(\cdot)$.
Inductive Transformation  In symbolic reasoning, a rule is considered as inductive if it is generalizable to different entities and times  (Xiong et al., 2022, 2023). For LLMs, this might be difficult to understand without enough training data since their fundamental principle is to learn the co-occurrence of tokens. To solve this problem, we introduce a new approach called inductive transformation. We globally map all the entity names in training data to some random names of the same type (e.g., "Knox Cunningham" $\rightarrow$ "Olivia Davis"; "Sydney" $\rightarrow$ "Cardiff"), and change the times based on a global time offset (e.g., the offset is "6 years": "June 1994" $\rightarrow$ "June 2000"; "2008" $\rightarrow$ "2014"). For each relation, we generate these random names with LLMs by providing several examples of existing entity names. To avoid confusing the model, we only use obscure names that do not exist in our dataset.
$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e)$ (6)
$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c))$
$a^{\prime\prime}=f_{T}(f_{E}(a))$
where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity name changing and time changing, respectively.
4 Experiments
We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation? (2) Can our two-step framework lead to better temporal reasoning performance? (3) Can LLMs benefit from our data on other reasoning tasks?
4.1 Experimental Setup
In addition to TempGraph QA, we consider the existing dataset, TimeQA  (Chen et al., 2021), which is based on the texts from Wikipedia. It contains two difficulty levels: The easy-level split tends to be information extraction while the hard-level split involves temporal reasoning, i.e., understanding the relation between different temporal expressions. We evaluate our framework on both datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider the accuracy based on perplexity for LLM evaluations. Compared with other LMs, LLMs are pre-trained on a wide variety of tasks, making them more generalizable at the cost of less conforming to the required format. This property actually hurts their performance under the commonly used metrics such as EM and F1. To solve this problem, we select the answers from a candidate set based on their perplexity from the LLM. For questions with multiple correct answers, we follow the strategy proposed in  (Chen et al., 2021) to get the best F1, EM and Acc among them. For TempGraph QA, we first calculate the metrics for each category, and then use the average as final scores to mitigate question category imbalance.
We compare our framework with leading LLMs including Llama2, GPT-3.5 and GPT-4. Llama2 Touvron et al. (2023b) is one of the most accepted open-source LLMs, while GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023) are two powerful closed-source models. We run Llama2 series on local machines (4-bit quantization is adopted for 70B version due to computational resource limitation), and use APIs provided by OpenAI for GPT series. We evaluate their few-shot ICL performance on the test set. To avoid LLMs generating free text which affects the evaluation with these metrics  (Yang et al., 2023a), we provide candidate answers for all questions under the ICL setting. The model versions and prompt templates are shown in Appendix B and C, respectively.
4.2 Implementation Details
We use Llama2-13B as the baseline due to computation resource limit. In experiments, we proved that our framework has comparable or even better temporal reasoning performance than one of the most powerful LLMs, GPT-4. We inject two adapters with selectors into the Llama2 model for the text-to-tempgraph translation and symbolic temporal reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform symbolic temporal reasoning on it, i.e., the adapters are used in sequence. In addition, we use GPT-3.5 for the story creation & verification in TempGraph QA, entity and relation extraction, temporal info identification, graph construction & verification on other datasets, and symbolic CoTs bootstrapping. These examples used in ICL are created by GPT-4 with manual verification. We choose these models due to the high quality of their generation. All the prompt templates are shown in Appendix C.
4.3 Main Results
Can our strategies guide LLMs to perform more reliable symbolic reasoning?   Figure 2 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TempGraph QA. To obtain a fair comparison, all the reasoning happens on the estimated graphs obtained from our text-to-graph translation module. It can be seen that LLMs learn better temporal reasoning with SFT than ICL. By providing CoTs with symbolic reasoning, the model performance gets further enhanced. There is an additional performance boost when using our data augmentation strategies. Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies. Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error). We measure the level of agreement among the three evaluators, and show the error rates from different types in Table 4. It can be seen that LLMs with ICL easily use either wrong information or erroneous external knowledge in reasoning. This issue can be alleviated with SFT, and we shows that using symbolic reasoning is better than free-text CoTs. Our data augmentation strategies further reduce all types of errors, and provides a promising direction for LLMs on reasoning tasks.
Figure 2: Performance comparison between different strategies of CoT generation on TempGraph QA.
Strategy ER-T1 ER-T2 ER-T3 ER-T4
ICL 0.13 0.13 0.31 0.10
SFT 0.04 0.17 0.13 0.10
SFT + bs 0.06 0.13 0.03 0.08
SFT + bs + aug 0.03 0.05 0.04 0.05
Table 4: Human evaluation on the CoTs generated by different strategies. T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error.
Model TempGraph QA TimeQA
Easy-mode Hard-mode
EM F1 Acc EM F1 Acc EM F1 Acc
GPT-3.5 (ICL-SP) 0.559 0.699 - 0.640 0.668 - 0.512 0.506 -
GPT-3.5 (ICL-CoT) 0.698 0.788 - 0.565 0.603 - 0.436 0.464 -
GPT-4 (ICL-SP) 0.704 0.818 - 0.716 0.742 - 0.571 0.546 -
$*{}^{*}$ (ICL-CoT) 0.798 0.833 - 0.662 0.693 - 0.618 0.636 -
Llama2-7B (ICL-SP) 0.402 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354
Llama2-7B (ICL-CoT) 0.548 0.686 0.578 0.367 0.425 0.393 0.302 0.354 0.360
Llama2-13B (ICL-SP) 0.423 0.609 0.526 0.439 0.493 0.450 0.427 0.481 0.437
Llama2-13B (ICL-CoT) 0.626 0.762 0.668 0.518 0.572 0.535 0.434 0.490 0.480
Llama2-70B (ICL-SP) 0.578 0.736 0.665 0.583 0.627 0.631 0.493 0.537 0.551
Llama2-70B (ICL-CoT) 0.760 0.838 0.796 0.552 0.612 0.623 0.447 0.501 0.512
Llama2-13B (SFT-SP) 0.550 0.720 0.652 0.412 0.449 0.455 0.362 0.404 0.412
Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379
Llama2-13B (TempGraph) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649
Table 5: Main results using different models and strategies. We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc). Results with * are evaluated on 1000 random test samples.
Figure 3: Performance comparison between different strategies on TimeQA. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies.
Can our two-step framework lead to better temporal reasoning performance?   Table 5 shows the comparison between different models on both TempGraph QA and TimeQA. First, we observe that among all the LLMs with ICL, GPT-4 has the best temporal reasoning capability as expected. For Llama2 series, larger model has better performance due to advanced context understanding and improved generalization. We also find that CoTs not always bring better temporal reasoning. In both easy-mode and hard mode of TimeQA, CoTs hurt larger LMs such as GPT-3.5, GPT-4 and Llama2-70B due to unreliable intermediate results and hallucinations. More importantly, our model, which is based on Llama2-13B, shows a comparable or even better temporal reasoning performance than GPT-4 on both datasets. To prove the effectiveness of our framework, we also provide the results of some alternative strategies (ICL, SFT-SP, SFT-CoT), where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively. Compared with our framework, there are no text-to-graph translation and symbolic reasoning involved in SFT-SP and SFT-CoT. It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets. Compared with original text, we actually provide an easier path from temporal reasoning questions to their answers.
Can LLMs benefit from our data on other reasoning tasks? Figure 3 shows the performance comparison between different strategies on TimeQA (ICL-SP, ICL-CoT, pre-training on TempGraph QA, SFT on TimeQA). It can be seen that the performance of in-context learning (either with standard I/O prompt or CoTs) is much lower than pre-training and SFT. In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning. More importantly, we observe that per-training on synthetic data improves the model performance compared with ICL, which proves the rationality of design and construction of our dataset. Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions. Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the reasoning capability of LLMs.
4.4 Ablation Study
We ablate different modules to see their contributions to the model performance. Table 6 shows the performance comparison between different configurations where TG, EK, SR denote temporal graph, external knowledge and symbolic reasoning, respectively. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. We use supervised fine-tuning on CoTs (SFT-CoT) as baseline. We improve the model performance 2.9% in EM, 3.2% in F1 and 0.6% in Acc by adding TG. Then we show that bootstrapping CoTs with symbolic reasoning brings further enhancement of 4.8% in EM, 4.3% in F1 and 2.5% in Acc. These is a significant boost in EM (5.9%), F1 (4.1%) and Acc (5.9%), when using our data augmentation strategies. Finally, the introduction of external knowledge exhibits meaningful results: it increases EM by 3%/1.5%, F1 by 0.6%/1.2%, and Acc by 4%/2.4% without/with data augmentation. From the ablation study, we give some insights for LLMs: (1) LLMs can benefit from (temporal) graph which is intuitive, concise and structured. (2) Given a reliable graph, symbolic reasoning brings better performance in reasoning than free-text CoTs. (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning. (4) The introduction of external knowledge such as mathematics and commonsense can augment the generation of LLMs.
Config EM F1 Acc
Baseline (SFT-CoT) 0.646 0.722 0.705
TG 0.675 0.754 0.711
TG + SR(bs) 0.723 0.797 0.736
TG + SR(bs + aug) 0.782 0.838 0.795
TG + EK + SR(bs) 0.753 0.803 0.776
TG + EK + SR(bs + aug) 0.797 0.850 0.819
Table 6: Ablation Study Results on TempGraph QA. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations.
5 Related Work
Text-based Temporal Reasoning   Recently, text-based temporal reasoning has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023). The vision here is to help LMs understand temporal concepts, which are fundamental elements in human knowledge. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust temporal reasoning.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge or extra temporal constraints. Although these methods make some progress, the underlying structure and logic of temporal reasoning are either ignored or not explicitly involved. Since LMs perform pre-training on human knowledge, the manual perspective of modelling might effectively improve their performance on this task.
Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and their capability of reasoning (Valmeekam et al., 2022; Wei et al., 2022; Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modelling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing the final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that requires reasoning. To the best of our knowledge, the exploration of the full capabilities of LLMs on temporal reasoning is still missing. Compared with other reasoning types, it is a form of integrated task that requires the knowledge of mathematics, logic and commonsense.
Neuro-symbolic Reasoning   Recent research shows advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Susskind et al., 2021; Hamilton et al., 2022). Knowledge graph is an important representation as the foundation of symbolic reasoning. Existing works on symbolic reasoning over KGs include  (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021). On top of that,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between text-based and temporal KG-based temporal symbolic reasoning. This connection brings the potential to extend these KG-based methods on text-based tasks. Some examples include  (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023). Different from these methods, which limit their application to certain datasets, our framework has better generalization due to the text-to-graph translation part.
6 Conclusion
TempGraph-LLM, a new paradigm for language models, has been proposed to improve their performance on temporal reasoning. Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning process. Experiments indicate that TempGraph-LLM achieves better performance than existing methods. An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning. Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks with good problem formalization and methodology extension.
Appendix A Dataset Statistics
In this section, we show statistics for our TempGraph QA dataset. Figure 4 shows the number of facts in training and test set, which determine the complexity of task to some extent. Figure 5 shows the types of questions in training and test set. Different types of questions are exemplified in Table 1. Among all the types, question 1 has the largest portion since it has five variants. Similarly, question 7 with two variants also has a larger portion. To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores.
Figure 4: Number of facts in TempGraph QA.
Figure 5: Types of questions in TempGraph QA.
Appendix B Model Versions
All the model versions used in our experiments are listed below:

### 20 Feb 2024 v2

@@ 3.2.2 Data Augmentation
Compared with conventional tasks, reasoning suffers more from data insufficiency since more information, such as evidence, arguments, and logics, are involved in intermediate steps  (Huang and Chang, 2023). To address the issue, we propose special strategies for data augmentation.
Essentially, our framework involves two steps: text-to-TG translation and symbolic temporal reasoning. At the test stage, reasoning happens on the noisy estimated graphs, which impacts the overall performance. Specifically, for the reasoning module, we train with ground-truth/verified graphs while in the test time, the estimated graph is used for symbolic reasoning. To improve the robustness of reasoning, we introduce some disturbances on the graphs during training. Note that the type of disturbances should be carefully designed in order to avoid confusing the model. We first investigate two types of disturbances: (i) remove irrelevant edges and (ii) replace edges by changing relation names, given in Eq. 5 and Eq. 6, respectively. An edge is considered irrelevant if it is not involved in both QA and CoT.
@@
Figure 3: We have several strategies for data augmentation: remove irrelevant edges, use relation synonyms and change entities/times.
where $P_{0}:=P(c|g,e,q)$ denotes the original conditional probability, irrelevant event $F_{irr}$ will be randomly removed from $g$, and disturbed event $F^{\prime}$ is obtained from the global mapping of relation names $f_{R}(\cdot)$.
Furthermore, to make sure LLM learns the underlying principle of symbolic temporal reasoning instead of just memorizing the training data, we introduce another two types of disturbances: (iii) globally map all the entity names in training data to some random names of the same type , and (iv) change the times based on a global offset, given in Eq. 7. For each relation, we generate these random names with GPT-3.5 by providing several examples of existing names.
$g^{\prime\prime}=f_{T}(f_{E}(g)),\ e^{\prime\prime}=f_{T}(e),$ (7)
$q^{\prime\prime}=f_{T}(f_{E}(q)),\ c^{\prime\prime}=f_{T}(f_{E}(c)),$
$a^{\prime\prime}=f_{T}(f_{E}(a))$
where $f_{E}(\cdot),f_{T}(\cdot)$ denote the global mapping of entity names and time changing, respectively.
4 Experiments
We aim to answer the following research questions in our experiments: (1) Can we guide LLMs to perform more reliable symbolic reasoning with the strategies of bootstrapping CoTs and data augmentation? (2) Can our two-step framework lead to better temporal reasoning performance? (3) Does these learned capabilities of temporal reasoning generalize to other tasks?
4.1 Experimental Setup
In addition to TGQA, we consider the two existing datasets, TimeQA  (Chen et al., 2021) and TempReason  (Tan et al., 2023a), which are constructed using Wikipedia articles, excerpts, or summaries. Compared with these datasets, TGQA has a better quality due to the controlled generation, and more diverse question categories (see examples from other datasets in Appendix B). We evaluate our framework on all the datasets with the metrics of token-level F1, exact match (EM) and perplexity-based accuracy (Acc). Besides choosing F1 and EM, which are two basic metrics for span-based QA tasks, we consider Acc for LLM evaluations, i.e., selecting from a candidate set the answer with the lowest perplexity from LLMs as their prediction. (See rationales for evaluation metrics selection in Appendix B)
We compare our framework [^1] Code and data available at https://github.com/xiongsi heng/TG-LLM. with leading LLMs including Llama2 Touvron et al. (2023a), GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023). We run the Llama2 family on local machines (4-bit quantization for Llama2-70B due to limited computational resources), and use APIs provided by OpenAI for GPT models. We evaluate their few-shot ICL performance on the test set. To avoid LLMs generating free text which affects the evaluation, we provide candidate answers for all questions. The model versions and prompt templates are provided in Appendix B and C, respectively. We also include the T5-based models  (Tan et al., 2023a; Yang et al., 2023a) for a comprehensive comparison. Results on TimeQA and TempReason reported in the original papers are used, and these models are fine-tuned and evaluated on TGQA.
4.2 Implementation Details
We use Llama2-13B as the baseline due to limited computational resources. We inject two adapters with selectors into the model for the text-to-TG translation and symbolic temporal reasoning. The adapters are trained in parallel. For inference, we first translate the original story into a temporal graph, and then perform symbolic reasoning on it, i.e., the adapters are used in sequence. In addition, we use GPT-3.5/GPT-4 for all the generation of story, graph and CoTs due to their high quality generation. All the prompt templates are given in Appendix C.
Figure 4: Performance comparison between different strategies of CoT generation on TGQA.
Strategy ER-T1 ER-T2 ER-T3 ER-T4
ICL 0.13 0.13 0.31 0.10
SFT 0.04 0.17 0.13 0.10
SFT + bs 0.06 0.13 0.03 0.08
SFT + bs + aug 0.03 0.05 0.04 0.05
Table 3: Human evaluation on the CoTs generated by different strategies. T1: using wrong info; T2: logical inconsistency; T3: external knowledge error; T4: temporal graph error. (See Appendix B for error type explanations)
  Model TGQA TimeQA TempReason
Easy-mode Hard-mode OBQA-L2 OBQA-L3
EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc EM F1 Acc
${}^{{\dagger}}$ 0.410 0.608 - 0.600 0.682 - 0.556 0.641 - 0.260 0.450 - 0.238 0.418 -
${}^{{\dagger}}$ 0.548 0.713 - 0.631 0.716 - 0.595 0.681 - 0.327 0.509 - 0.288 0.468 -
${}^{{\dagger}}$ 0.640 0.778 - - - - - - - 0.318 0.496 - 0.261 0.430 -
${}^{{\dagger}}$ 0.435 0.633 - 0.614 0.704 - 0.582 0.673 - 0.336 0.516 - 0.285 0.449 -
${}^{{\dagger}}$ 0.461 0.660 - 0.637 0.723 - 0.605 0.693 - 0.374 0.549 - 0.334 0.493 -
GPT-3.5 (ICL-SP) 0.598 0.699 - 0.640 0.668 - 0.512 0.506 - 0.303 0.409 - 0.365 0.453 -
GPT-3.5 (ICL-CoT) 0.706 0.788 - 0.565 0.603 - 0.436 0.464 - 0.340 0.478 - 0.243 0.348 -
$*{}^{*}$ (ICL-SP) 0.772 0.829 - 0.716 0.742 - 0.571 0.546 - 0.454 0.525 - 0.431 0.485 -
$*{}^{*}$ (ICL-CoT) 0.821 0.865 - 0.662 0.693 - 0.618 0.636 - 0.388 0.480 - 0.352 0.447 -
Llama2-7B (ICL-SP) 0.415 0.596 0.447 0.352 0.408 0.367 0.341 0.404 0.354 0.205 0.344 0.226 0.044 0.084 0.153
Llama2-7B (ICL-CoT) 0.548 0.686 0.578 0.367 0.425 0.393 0.302 0.354 0.360 0.233 0.410 0.263 0.179 0.357 0.187
Llama2-13B (ICL-SP) 0.440 0.609 0.526 0.439 0.493 0.450 0.427 0.481 0.437 0.284 0.452 0.289 0.189 0.370 0.183
Llama2-13B (ICL-CoT) 0.628 0.762 0.668 0.518 0.572 0.535 0.434 0.490 0.480 0.330 0.498 0.368 0.242 0.391 0.272
Llama2-70B (ICL-SP) 0.618 0.736 0.665 0.583 0.627 0.631 0.493 0.537 0.551 0.358 0.491 0.387 0.128 0.181 0.148
Llama2-70B (ICL-CoT) 0.761 0.838 0.796 0.552 0.612 0.623 0.447 0.501 0.512 0.325 0.477 0.345 0.303 0.393 0.318
Llama2-13B (SFT-SP) 0.550 0.720 0.652 0.412 0.449 0.455 0.362 0.404 0.412 0.337 0.506 0.338 0.244 0.408 0.253
Llama2-13B (SFT-CoT) 0.646 0.722 0.705 0.550 0.586 0.564 0.332 0.391 0.379 0.256 0.433 0.281 0.285 0.409 0.305
Llama2-13B (SFT-TGSR) 0.797 0.850 0.819 0.664 0.691 0.673 0.631 0.664 0.649 0.424 0.522 0.432 0.356 0.469 0.399
Table 4: Main results using different models and strategies. We report exact match (EM), token-level F1 scores, and perplexity-based accuracy (Acc). Note: (1) Results with ${}^{{\dagger}}$ are reported in the original papers. We fine-tune and evaluate the models on our dataset. (2) Results with * are evaluated on 1000 random test samples.
Figure 5: Performance comparison between different strategies on TimeQA and TempReason. To obtain a fair comparison, we use Llama2-13B as the base model for all strategies. The basic strategy is in-context learning with standard Input/Output (ICL-SP), which is used to calculate the performance changes.
4.3 Main Results
Can our strategy bring more reliable symbolic reasoning?   Figure 4 shows the comparison between ICL with CoTs, SFT with CoTs, bootstrapping CoTs, and data augmentation on TGQA. It can be seen that LLMs learn better temporal reasoning with SFT than ICL. By providing CoTs with bootstrapping and our data augmentation strategies, the model performance gets further enhanced. Furthermore, inspired by  (Wang et al., 2023), we manually check 100 CoTs generated by different strategies. Evaluators are asked to classify the errors into four types (using wrong info, logical inconsistency, external knowledge error, and temporal graph error). It can be seen in Table 3 that our strategies reduce all types of error rates (ER).
Can our two-step framework lead to better temporal reasoning performance?   Table 4 shows the comparison between different models on all datasets. First, we observed that among all the LLMs with ICL, GPT-4 has the strongest performance as expected. For the Llama2 family, larger models have better performance due to advanced context understanding and improved generalization. We also found that CoTs not always bring better temporal reasoning due to unreliable intermediate results and hallucinations. More importantly, our model (SFT-TGSR), which is based on Llama2-13B, shows a comparable or even better performance than GPT-4 on all datasets. To prove the effectiveness of our framework, we also provide the results of some alternative strategies, where SFT-SP and SFT-CoT denote supervised fine-tuning on standard Input/Output prompts and CoTs, respectively. It can be seen that the two-step framework brings a substantial improvement in model performance on different datasets. We hypothesize that this improvement in performance is because our two-step neuro-symbolic process provides an easier path toward answering the temporal questions for LLMs.
Does these learned capabilities of temporal reasoning generalize to other tasks?  Figure 5 shows the comparison between different strategies on TimeQA and TempReason. In our framework, LLMs can learn the capability of text-to-graph translation and symbolic reasoning, which brings better temporal reasoning. More importantly, we observe that per-training on TGQA improves the model performance compared with ICL. Also, it can be concluded that these necessary capabilities in temporal reasoning are generalizable to different data distributions. Since the synthetic dataset is fully controllable and requires minimal supervision, we actually provide a way of improving the temporal reasoning capability of LLMs.
4.4 Ablation Study
We ablate different modules to see their contributions to the performance. Table 5 shows the performance comparison between different configurations. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. From the ablation study, we obtain some insights: (1) LLMs can benefit from explicitly presented (temporal) graph which is intuitive, concise and structured. (2) Given a reliable graph, symbolic reasoning brings better performance than free-text CoTs. (3) Data augmentation is necessary for LLMs to perform complex tasks such as temporal reasoning. (4) The introduction of external knowledge such as mathematics and commonsense can augment the generation.
5 Related Work
Language-based Temporal Reasoning  Recently, language-based TR has gained substantial research focus  (Nylund et al., 2023; Hu et al., 2023; Gurnee and Tegmark, 2023; Liu et al., 2023). The vision here is to help LMs understand temporal concepts. Existing methods mainly solve this problem with time-aware modeling. For example,  (Rosin et al., 2022; Pereira, 2022; Tan et al., 2023a) propose pre-training/fine-tuning strategies for robust TR.  (Ning et al., 2019; Zhou et al., 2020; Yang et al., 2020; Zhou et al., 2021; Yang et al., 2023a) design auxiliary objectives to introduce external knowledge. Although these methods make some progress, the underlying structure and logic of TR are either ignored or not explicitly involved.
Reasoning towards LMs   Although LLMs exhibit some emergent behaviors, it is still unknown whether they can actually perform reasoning and how strong their capability of reasoning is (Huang and Chang, 2023). Existing methods that try to elicit or improve reasoning can be divided into two directions: reasoning-involved modeling or hybrid methods. For example,  (Wei et al., 2022; Kojima et al., 2022; Yao et al., 2023) use prompt engineering where some examples including intermediate reasoning steps are provided in context.  (Talmor et al., 2020; Nye et al., 2021; Wang et al., 2023; Ho et al., 2023) fine-tune LMs with these information before producing final answers. On the other hand,  (Li et al., 2023; Kynoch et al., 2023; Zhuang et al., 2023; Xu et al., 2023) combine LMs with domain-specific tools, empowering the model to perform more complex tasks that require reasoning.
Neuro-symbolic Reasoning Recent research has shown the advantages of combining neural networks with symbolic reasoning  (Sarker et al., 2021; Hamilton et al., 2022). In particular, knowledge graphs (KGs) as the foundation of (temporal) symbolic reasoning has been widely adopted in the past (Yang et al., 2017; Payani and Fekri, 2019a, b; Yang and Song, 2020; Qu et al., 2021). Moreover,  (Liu et al., 2022; Xiong et al., 2022, 2023) develop approaches for symbolic reasoning over temporal KGs. In this paper, we build the connection between language-based and temporal KG-based temporal symbolic reasoning. This connection brings the potential for extending these KG-based methods to language-based tasks. Different from existing methods (Luo et al., 2023; Zhang et al., 2023; Yuan et al., 2023), which limit their application to certain datasets, our framework offers enhanced generalization and usability, largely attributed to its innovative use of text-to-graph translation as a precursor to symbolic reasoning.
Config. EM F1 Acc
Baseline (SFT-CoT) 0.646 0.722 0.705
TG 0.675 0.754 0.711
TG + SR(bs) 0.723 0.797 0.736
TG + SR(bs + aug) 0.782 0.838 0.795
TG + EK + SR(bs) 0.753 0.803 0.776
TG + EK + SR(bs + aug) 0.797 0.850 0.819
Table 5: Ablation study results on TGQA. To obtain a fair comparison, we use Llama2-13B as the base model for all configurations. TG: temporal graph, SR: symbolic reasoning, EK: external knwoledge, bs: bootstrapping, aug: data augmentation.
6 Conclusion
TG-LLM, a new framework for language models, has been proposed to improve their performance on temporal reasoning. Besides producing the final answers, our framework also provides the temporal graph and symbolic reasoning steps. Experiments indicate that TG-LLM achieves better performance than existing methods. An interesting direction for future work is to extend it to more complex applications such as induction and abductive reasoning. Due to the graph structure and capability of symbolic reasoning, it is promising to improve the performance of LLMs on these tasks.
Limitations
Graph-augmented approaches such as TG-LLM help language models better learn related concepts from the perspective of graph. Although we demonstrate TG-LLM has good performance on understanding temporal relations, it still needs adaptations for temporal commonsense reasoning  (Zhou et al., 2019; Qin et al., 2021). Explicit in-context integration of commonsense presents opportunities for this task. Further, we mainly improve the capability of LLMs by introducing a new paradigm and providing more plausible and informative training data. There can be opportunities such as simulating an environment to provide feedback to LLMs  (Yao et al., 2023). For example, we can verify the generated TG based on prior knowledge such as the time gap between someone’s birth date and death date, i.e., a person’s lifespan, should fall into a certain range. In this way, we can further improve the quality of generation.
Ethics Statement
In this paper, we adopt YAGO11k for fine-tuning the language models. The dataset is publicly available, and is for research purposes only. We also use GPT model to generate text based on YAGO11k, for which OpenAI has been committed to addressing ethical considerations. In addition, we adopt TimeQA and TempReason for evaluation. Both datasets are publicly available, and are for research purposes only. However, they may still contain improper or harmful content. None of such content reflects the opinions of the authors.
Appendix A TGQA Dataset Statistics
In this section, we show statistics for our TGQA dataset. Figure 6 shows the number of facts distribution in training and test set, which determine the complexity of the task to some extent. Figure 7 shows the types of questions distribution in training and test set, where Q0: "Which event occurred first, <Event_A> or <Event_B>?", Q1: "Given the following <N> events: <Event_A>, <Event_B>, <Event_C>, <Event_D>, $\cdots$, which event is the first/second/third/fourth/$\cdots$ one in the chronological order?", Q2: "How long did the event <Event_A> last?", Q3: "", Q4: "How much time passed between the start of <Event_A> and the start of <Event_B>?", Q5: "What happened right before/after <Event_A> started?", Q6: "When did the <Event_A> occur?", Q7: "True or false: <Event_A> and <Event_B> happened at the same year?", Q8: "True or false: <Event_A> was still happening when <Event_B> started?". Note that we give eight categories of question in Table 2, since in a strict sense Q0 and Q1 can be considered as the same type. Among all the question types, Q1 has the largest portion since it has multiple variants. Similarly, Q7 with two variants also has a larger portion. To mitigate question category imbalance, we first calculate the metrics for each category, and then use the average as final scores. We also provide examples for the global mapping for entity names (Table 6) and external knowledge (Table 7) used in our dataset.
Entity Name Mapping
Ori New
Beverly_Adams Isabella_Thompson
Edmonton Lancaster
Al_Gore Chris_Evans
Carrara_Stadium Maplewood_Arena
Bend_Sinister_(novel) Lakefield_Chronicles_(novel)
$\cdots$ $\cdots$
Table 6: We use a global mapping for entity names from GPT-3.5 to avoid corpus leakage.
External Knowledge
1885 before 1893 before 1916 before 1918 before 1922
before 1928 before 1941
1918 - 1916 = 2
1928 - 1893 = 35
1928 - 1922 = 6
1941 - 1918 = 23
2 < 6 < 23 < 35
Table 7: We integrate the necessary mathematics in context as external knowledge for temporal reasoning.
Appendix B Experiment Details
In this section, we present more experiment details for further research. We include data augmentation examples, model versions and evaluation tasks & metrics.
Temporal Graph:
@@ Appendix B Experiment Details
Q: True or false: event (John Thompson owned Pearl
@@
Graph-based QAs:
@@
Q: True or false: event (James Brown owned Iris Inn)
@@
Table 8: We propose special data augmentation strategies for symbolic reasoning over (temporal) graphs.
Data Augmentation Examples  We show with an example our data augmentation strategies in Table 8. For TGs with relation synonym replacement or irrelevant edges removal, there is no need for change on Graph-based QAs. For TGs with entity names/times mapping, we also need to change with the corresponding entity names/times in Graph-based QAs to ensure the consistency.
Model Versions  All the model versions of LLMs used in our experiments are listed below: